OpenStudy (elleblythe):

How to solve for cot147 MANUALLY? How do I get the values of 147 if it isn't a special angle on the circle?

OpenStudy (anonymous):

suck my aaass

OpenStudy (anonymous):

\[-((1/8 (\sqrt(5)-1)+1/4 \sqrt(3/2 (5+\sqrt(5))))/\sqrt(2)-(1/8 \sqrt(3) (\sqrt(5)-1)-1/4 \sqrt(1/2 (5+\sqrt(5))))/\sqrt(2))/((1/8 \sqrt(3) (\sqrt(5)-1)-1/4 \sqrt(1/2 (5+\sqrt(5))))/\sqrt(2)+(1/8 (\sqrt(5)-1)+1/4 \sqrt(3/2 (5+\sqrt(5))))/\sqrt(2))\]

OpenStudy (elleblythe):

@bakur sorry but I don't really understand your solution...??